OpenStudy (anonymous):

Solve the equations:

OpenStudy (anonymous):

\[4x^2-5xy+y^2=10, 3x^2-3xy-y^2=5\]

OpenStudy (anonymous):

^system of equations

OpenStudy (anonymous):

anyone there ?? :/

OpenStudy (anonymous):

so, what i would do is this: 4x^2−5xy+y^2=10 -(3x^2−3xy−y^2=5 x^2-2xy=5

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

then plug in the values and solve for x

OpenStudy (anonymous):

i multiplied eqn 1 by 3 and 2 by 5 and then subtracted both i got -3x^2+8y^2=5

OpenStudy (anonymous):

am i doing it right?

OpenStudy (anonymous):

is there any online system of equations solver?

OpenStudy (raden):

wolfram :)

OpenStudy (raden):

actually, this question can solved by elimination and subtitution method

OpenStudy (anonymous):

how plz tell me??

OpenStudy (anonymous):

plz can someone tell me what is the type of these equations??

OpenStudy (anonymous):

i've asked this question before ;)

OpenStudy (anonymous):

the answer is correct !! but i should be able to see the steps

OpenStudy (anonymous):

thanks btw

OpenStudy (anonymous):

i found a question like this in my book

OpenStudy (anonymous):

it is an example... which is solved!!

OpenStudy (raden):

well, there is a nice way to solve it! 4x^2 - 5xy + y^2 = 10 ... (1) 3x^2 - 3xy - y^2 = 5 ... (2) division from (1) and (2), giving us : (4x^2 - 5xy + y^2)/(3x^2 - 3xy - y^2) = 10/5 (4x^2 - 5xy + y^2)/(3x^2 - 3xy - y^2) = 2 cross product : 4x^2 - 5xy + y^2 = 6x^2 - 6xy - 2y^2 simplify ! 2x^2 - xy - 3y^2 = 0 factor out ! (2x - 3y)(x + y) = 0 the solution of this, are x = 3y/2 or x = -y

OpenStudy (raden):

now, subtitute of both (x's) to one of equations above! example, i take the 1st equation : 4x^2−5xy+y^2=10 try for the first x, is x=3y/2 4(3y/2)^2−5(3y/2)y+y^2=10 4(9/4)y^2 - 15y/2 + y^2 = 10 simplify ! y^2 = 4 y = +- 2 subtitute back to x = 3y/2 so, x1 = 3(2)/2 = 3 ----> solution (3,2) x2 = 3(-2)/2 = -3 -----> solution (-3,-2)

OpenStudy (raden):

now, for u the rest :)

OpenStudy (raden):

take x = -y and subtitute to 4x^2−5xy+y^2=10, so that all terms in y solve for y, then subtitute back to x=-y to get the value of x u'll get order pairs as solutions